OpenStudy (anonymous):

Write an equation in point-slope form of the line through the point (–1, 2) and with slope m = 2.

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) (x1,y1) is a coordinate and m is the slope, just plug those in

OpenStudy (anonymous):

im like lost on slope & im taking a test on it right now :(

OpenStudy (anonymous):

You were given the slope and a point, all you have to do is put them in the right spot. y-2=2(x+1)

OpenStudy (anonymous):

thanks so much